• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kai

    Log entry time 01:19:41 on November 22, 2009

    Entry number 301695

    keyword=owl shift summary

    16:00 shift start with on beam, MCC said they have RF problem.

    16:00 We asked for controlled access. PVDIS guys went to the hall to add delayed cable on the pion path on the left arm, and replace a scaler on the right arm. Compton guys also went to the hall to add some shielding to lower the low energy phone background.

    21:00 MCC called they fixed the RF problem. However, they happened to find there is unterminated valve in the linarc which is problematic, and they said they are not able to deliver the beam until mid night.

    22:00 change from controlled access to power permit

    24:00 MCC is ready to send beam, and we are following the procedure to recover the beam.